OpenStudy (winsteria):

WILL MEDAL AND FAN The base of a triangle is 5/8 ft. The height is 1/3 ft. How much greater is the base of the triangle than the height? Express your answer in lowest terms.

OpenStudy (winsteria):

@abswag13 @StudyGurl14

OpenStudy (studygurl14):

Do this: \(\large\frac{5}{8}\div\frac{1}{3}\rightarrow\frac{5}{8}\times\frac{3}{1}\rightarrow\frac{15}{8}\)

OpenStudy (anonymous):

that a sluty picture

OpenStudy (winsteria):

@KendrickLamar2014

OpenStudy (kendricklamar2014):

@stupidk1234 thats not nice, stop or ill report u

OpenStudy (kendricklamar2014):

Give me a moment @winsteria

OpenStudy (winsteria):

it's not even me though lol

OpenStudy (winsteria):

@abswag13

OpenStudy (kendricklamar2014):

Ok back

OpenStudy (anonymous):

sorry im busy now

OpenStudy (kendricklamar2014):

5/8 - 1/3 = ? @winsteria

OpenStudy (anonymous):

i wan to help but i cant

OpenStudy (winsteria):

don't you have to find a common denominator?

OpenStudy (anonymous):

i thought that was your picture to

OpenStudy (winsteria):

it's okay @abswag13

OpenStudy (kendricklamar2014):

No

OpenStudy (winsteria):

and no it's Sky Ferreira

OpenStudy (winsteria):

7/24? @KendrickLamar2014

OpenStudy (kendricklamar2014):

Correct

OpenStudy (kendricklamar2014):

and that is ur answer

OpenStudy (winsteria):

Uh can't that be simplified @KendrickLamar2014

OpenStudy (kendricklamar2014):

No

OpenStudy (winsteria):

@Alfa_female @Light&Happiness

OpenStudy (kendricklamar2014):

Do u not belif me @winsteria ??

OpenStudy (winsteria):

idk I just want to make sure others agree so I don't get it wrong

OpenStudy (anonymous):

Is there multi choice?

OpenStudy (winsteria):

No

OpenStudy (anonymous):

I double checked, \[\frac{ 7 }{ 24 }\] is correct

OpenStudy (winsteria):

okay thank youuu both

OpenStudy (kendricklamar2014):

You'r Welcome

OpenStudy (anonymous):

7/24

OpenStudy (anonymous):